OpenStudy (anonymous):

Find f(x+3) for the following function. Please show all of your work. f(x) = x^2-7x+7 Find f(x+3) for the following function. Please show all of your work. f(x) = x^2-7x+7 @Mathematics

jimthompson5910 (jim_thompson5910):

f(x) = x^2-7x+7 f(x+3) = (x+3)^2-7(x+3)+7 ... replace EVERY x with 'x+3' f(x+3) = x^2+6x+9-7(x+3)+7 f(x+3) = x^2+6x+9-7x-21+7 f(x+3) = x^2-x-5

OpenStudy (anonymous):

f(x+3)= (x+3)^2 -7(x+3)+7 -> x^2-x-5

OpenStudy (anonymous):

so which answer is correct? These are two different ones. lol

jimthompson5910 (jim_thompson5910):

we both got the same answer, look at the last line of my work and after the "->" for the other poster btw answer is x^2-x-5 (which we both got)

OpenStudy (anonymous):

different way but same(answers)

OpenStudy (anonymous):

okay great. Thanks so much. I saw it differently because I dont understand this stuff. lol